find x.
FGH is equilateral with FG =x + 5, GH = 3x - 9, and FH = 2x -2.

Answer by actuary(112)

In an equilateral triangle, all sides are equal,
x+5=3x-9=2x-2
Take x+5=3x-9 and solve for x
14=2x
Solve for x
x=7
Take 3x-9=2x-2 and solve for x
x=7
So x=7
I hope that this helps.